a) assuming thatwater has density of exactly 1g/cm^3.
find the mass of one cubic meter of water in kg.
b) suppose it takes 10 hr to drain a container of 5700 m^3 of water, what is the "mass flow rate " in kg/sec of water from the container.
1
Expert's answer
2018-09-07T16:20:08-0400
a)
ρ=1 g/cm^3 =1 (〖10〗^(-3) kg)/(〖10〗^(-2) m)^3 =1 (〖10〗^(-3) kg)/(〖10〗^(-6) m^3 )=1000 kg/m^3
The mass of one cubic meter of water in kg:
m=ρV=1000 kg/m^3 (1m^3 )=1000 kg.
b) The "mass flow rate” in kg/sec of water from the container:
dm/dt=ρ dV/dt=1000 kg/m^3 ((5700 m^3)/10(3600)s)=158 kg/s.